OpenStudy (anonymous):

name 3 points that lie on the line y=3x-1

OpenStudy (lgbasallote):

are you familiar with the plotting of points method?

OpenStudy (anonymous):

Pick 3 values of x and calculate the y's to go with them....

OpenStudy (anonymous):

can you give me an example please?

OpenStudy (anonymous):

Pick a value of x...

OpenStudy (anonymous):

2

OpenStudy (anonymous):

y=3x-1 so what's y?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

Now do another two....

OpenStudy (anonymous):

y=3x-1 y=3(3)-1 y=9-1 y=8 y=3(5)-1 y=15-1 y=14

OpenStudy (anonymous):

what would i do for y=4 ?